OpenStudy (anonymous):

The function f(x) is continuous at 0. Show that f is differentiable at 0, and find f prime at 0. Hint: Look at the Newton quotient for f at 0 and use L’Hospital’s rule.

OpenStudy (anonymous):

f(x) = (sin x)/x when x is not = 0 = 1 when x = 0

OpenStudy (zarkon):

where are you stuck?

OpenStudy (anonymous):

not sure where to start

OpenStudy (zarkon):

look at \[\lim_{x\to 0}\frac{f(x)-f(0)}{x-0}\]

OpenStudy (anonymous):

use newtons quotient

OpenStudy (anonymous):

so after applying HR i get the limit as x->0 [xcos(x)-sin(x)]/x^2

OpenStudy (anonymous):

u used lhopitals? you sure \[\frac{f'(x)}{g'(x)}\]

OpenStudy (anonymous):

nvm i got it

OpenStudy (anonymous):

i got 0

OpenStudy (anonymous):

f'(x)= cos(x)/1 at 0 = 1

OpenStudy (zarkon):

0 is correct

OpenStudy (anonymous):

so that shows it's differentiable at 0?

OpenStudy (zarkon):

yes...if the limit did not exist then it would not be differentiable....the limit did exist so it is differentiable

OpenStudy (anonymous):

i'm so confused here are you guys talking about newtons quotient?

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

so now what is f prime at 0?

OpenStudy (zarkon):

0

OpenStudy (anonymous):

how?

OpenStudy (zarkon):

it is the limit you got

OpenStudy (anonymous):

oh i see thanks